If you weigh 660 N on the earth, what would be your weight on the surface of a neutron star that has the same mass as our sun and a diameter of 16.0 km ?

Take the mass of the sun to be ms = 1.99×1030 kg , the gravitational constant to be G = 6.67×10−11 N⋅m2/kg2 , and the acceleration due to gravity at the earth's surface to be g = 9.810 m/s2 . Express your weight wstar in newtons
Answer #1

On earth :

We = weight on earth = 660 N

m = mass

ge = acceleration due to gravity on earth = 9.810 m/s2

weight on earth is given as

We = m ge

660 = m (9.810)

m = 67.3 kg

On neutron star :

ms = mass of star = 1.99 x 1030 kg

R = radius = diameter/2 = 16/2 = 8 km = 8000 m

Weight on star is given as

Wstar = G ms m/R2

Wstar = (6.67 x 10-11) (1.99 x 1030) (67.3) /(8000)2

Wstar = 1.4 x 1014 N